OpenStudy (anonymous):

(4x-1)-(7x+6) simplify

OpenStudy (theviper):

\((4x-1)-(7x+6)\) \(=4x-1-7x-6\) \(=-3x-7\)

OpenStudy (anonymous):

why did you change the plus sigh to a minus?

OpenStudy (theviper):

Because it's - sign before brackets

OpenStudy (anonymous):

& in the final product, how did you get 7?

OpenStudy (theviper):

4x-7x=-3x & -1-6=-7